OpenStudy (anonymous):

WILL SOMEONE PLEASE ANSWER THIS! :) Solve for m in D m/v, if D= 5.1 and v = 0.3 the / is for a fraction.

OpenStudy (phi):

you can't solve unless you have an equation. You can replace the letters with numbers, and simplify 5.1m/0.3 can you divide 0.3 into 5.1?

OpenStudy (anonymous):

0.05

OpenStudy (phi):

it is the same problem as 51/3

OpenStudy (anonymous):

D = m / v D = 5.1 v = 0.3 5.1 = m / .3 5.1 * .3 = m m=153

OpenStudy (phi):

if it is D= m/v then m= 0.3* 5.1 = 1.53